40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 20 14:42:34 2011, steve, Update, SUS, oplev gains today 
    Reply  Wed Jul 20 19:24:12 2011, rana, Update, SUS, oplev gains today Untitled.pnga.png
       Reply  Wed Jul 20 19:48:03 2011, Jamie, Update, SUS, Re: oplev gains today 
       Reply  Thu Jul 21 09:04:59 2011, valera, steve, Update, SUS, oplev gains were not optimized 
       Reply  Mon Jul 25 00:35:44 2011, rana, Update, SUS, something wrong with ETMY LR sensor a.pngbad.png
          Reply  Tue Jul 26 11:38:45 2011, steve, Update, SUS, ETMY bias checked 
    Reply  Thu Jul 21 16:06:35 2011, steve, Update, SUS, oplev gains today checked by conlog 
Message ID: 5025     Entry time: Mon Jul 25 00:35:44 2011     In reply to: 5004     Reply to this: 5037
Author: rana 
Type: Update 
Category: SUS 
Subject: something wrong with ETMY LR sensor 

a.png

Looks like either the LR OSEM is totally mis adjusted in its holder or the whitening eletronics are broken.

Also looks like the ETMY is just not damped at 1 Hz? How can this be?

I look at the SUS_SUMMARY screen which apparently only Steve and I look at:

bad.png

Looks like the suspensions have factor of 10-100 different gains. Why?

**  The ETMY just doesn't behave correctly when I bias it. Both pitch and yaw seem to make it do yaw. I leave this for Jamie to debug in the morning.

***  Also, the BIAS buttons are still broken - the HOPR/LOPR limits ought to be 5000 and the default slider increment be 100. Also the YAW button readback doesn't correctly show the state of the BIAS.

****  And.....we have also lost the DAQ channels that used to be associated with the _IN1 of the SUSPOS/PIT/YAW filter modules. Please put them back; our templates don't work without them.

ELOG V3.1.3-